OpenStudy (anonymous):

What is x^2+4 divided by x-2

OpenStudy (lanhikari22):

Check if you can factor x^2 + 4, somehow.

OpenStudy (anonymous):

(x-2)(x-2)?

OpenStudy (lanhikari22):

FOIL that. You get: x^2 -2x - 2x + 4 = x^2 -4x + 4, so no.

OpenStudy (lanhikari22):

I feel like you have a problem in the question. Are you sure it's not x^2 - 4?

OpenStudy (anonymous):

yesI am sure..

OpenStudy (lolacole12):

x2+4x−2 =x3−2x+4x

OpenStudy (lanhikari22):

I guess you could complete the square. Or just do long division

OpenStudy (lolacole12):

Let's solve your equation step-by-step. x2+4x−2=x3−2x+4x x2+4x−2=x3+2x Step 1: Subtract x^3+2x from both sides. x2+4x−2−(x3+2x)=x3+2x−(x3+2x) −x3+x2+2x−2=0 Step 2: Factor left side of equation. (−x+1)(x2−2)=0 Step 3: Set factors equal to 0. −x+1=0 or x2−2=0 x=1

OpenStudy (lolacole12):

mY Brother had to do this

OpenStudy (anonymous):

Thank you.

OpenStudy (lolacole12):

Let's simplify step-by-step. x3−2x+4x =x3+−2x+4x Combine Like Terms: =x3+−2x+4x =(x3)+(−2x+4x) =x3+2x Answer: =x3+2x

OpenStudy (lolacole12):

My bro got 100% on test

OpenStudy (lanhikari22):

How is that related to simply dividing x^2+4 by x-2

OpenStudy (lolacole12):

x3−2x+4x

OpenStudy (lolacole12):

x2+4x−2 =x3−2x+4x

OpenStudy (lolacole12):

medal pls and bye now

OpenStudy (lanhikari22):

what? How is x^2 + 4x - 2 = x^3 - 2x + 4x? even if you multiply all by x you'd get x^3 + 4x^2 -2x